OpenStudy (anonymous):

Solve. 8 = 2x + 4 ? answers: x = -4 x = -1 x = 0 x = 7

OpenStudy (zepp):

2x = 8-4 2x = 4 x = 4/2 x = 2

OpenStudy (anonymous):

2x2 =4 and 4+4 =8 so it would be 4

OpenStudy (zepp):

There's either a mistake in the question, or the answers you copied are wrong.

OpenStudy (anonymous):

x=4

OpenStudy (anonymous):

srry i ment 2

OpenStudy (anonymous):

ya i got 2 also i think my answers are wrong unless it is -4?

OpenStudy (anonymous):

2

OpenStudy (zepp):

If x = -4 8 = 2x + 4 ? 8 = 2(-4) + 4 8 = -8 + 4 8 =/= -4

OpenStudy (anonymous):

thanks